# Reward-Forcing: Autoregressive Video Generation with Reward Feedback

### v1

@@ 3.2 Training with ODE and Reward Guidance
LTX-Video [16] 1.9B $768{\times}512$ 8.98 13.5 80.00 82.30 70.79
Wan2.1 [55] 1.3B $832{\times}480$ 0.78 103 84.26 85.30 80.09
@@
NOVA [11] 0.6B $768{\times}480$ 0.88 4.1 80.12 80.39 79.05
Pyramid Flow [24] 2B $640{\times}384$ 6.7 2.5 81.72 84.74 69.62
@@
SkyReels-V2 [6] 1.3B $960{\times}540$ 0.49 112 82.67 84.70 74.53
MAGI-1 [45] 4.5B $832{\times}480$ 0.19 282 79.18 82.04 67.74
CausVid [60] 1.3B $832{\times}480$ 17.0 0.69 81.20 84.05 69.80
Self Forcing [20] 1.3B $832{\times}480$ 17.0 0.69 84.31 85.07 81.28
@@
Optimization with Reward Feedback
After initializing the model with an ODE-based process that teaches motion synthesis, we introduce a reward-guided optimization stage to enhance video quality. Specifically, we adopt ImageReward [57] as our reward model and incorporate it into training in a differentiable manner to directly guide the video diffusion model.
Let the generated video be denoted as $\hat{x}_{1:T}=G_{\theta}(z)$, where $G_{\theta}$ is the autoregressive video generator parameterized by $\theta$, and $z$ is the latent input. The reward model $\mathcal{R}(\cdot)$ assigns a scalar reward indicating perceptual quality. We define the reward-guided objective as:
$\mathcal{L}_{\text{reward}}(\theta)=-\mathbb{E}_{z\sim\mathcal{Z}}\left[\mathcal{R}(\hat{x}_{T})\right].$ (7)
@@ 4 Experimental Results
4.1 Baseline Methods
We benchmark three classes of approaches: (i) standard bidirectional diffusion models, (ii) autoregressive methods that render videos one (latent) frame at a time, and (iii) autoregressive methods that synthesize videos in temporal chunks. For standard diffusion models, we include LTX-Video [16] and Wan2.1 [54] which is the base model used by [60, 20] and our method. For frame-wise autoregressive models, we include NOVA [11] which formulates the video generation problem as non-quantized autoregressive modeling of temporal frame-by-frame prediction and spatial set-by-set prediction, Pyramid Flow [24] that interprets the original denoising trajectory as a series of pyramid stages and Self Forcing [20] that trains the autoregressive models by relying on self-generated frames instead of ground-truth ones. For Chunk-wise models, we include SkyReels-V2 [6], MAGI-1 [45] that utilizes transformer based VAE and increased noise levels among chunks that are being generated, CausVid [60] that uses causal attention at training time and self-rollout at inference time and Self Forcing [20] which utilizes similar architectures as its frame-wise counterpart.
4.2 Implementation Details
Similar to [60, 20], our model is based on Wan2.1-T2V-1.3B [54]. The model is first distilled into a 4-step model using DMD. For ODE initialization, we also directly sample 1.4K trajectories from the bidirectional model. Our work uses Self-Rollout during training which is the same as SelfForcing [20]. For our main method, we only employ loss from the reward models. For our reward plus distillation setting, we train the model using DMD together with reward loss as shown in the previous section. Our method also works for both chunk-wise and frame-wise autoregressive generation. In our implementation, since we only need to sample the ODE trajectory from the teacher model, we don’t need any real training dataset. Thus our overall approach remains data free which is similar to that of Self Forcing [20]. Similar to previous works [20], we enable EMA [22] during the training process.
Computing Infrastructure
Hyperparameters
Here we list the hyperparameters we used for our experiment. We use a batch size of 8 after training with the ODE trajectory. For the optimizer, we use AdamW with $\beta_{1}=0,\beta_{2}=0.999,\epsilon=1e-8,weight\_decay=0.01$ with a learning rate of 2e62e-6. For EMA, we use a decay of 0.99. When combing distillation loss with image reward loss, we normalize the losses so that they are on the same scale. When comparing with baseline methods such as DMD or Causvid, we use their author-provided checkpoints.
Note that our training process eliminates the second distillation stage used in CausVid [60] and Self Forcing [20]. As a result, there is no need to load the teacher and critic models or train the critic model to approximate the generator distribution, making the training process after ODE initialization significantly more lightweight and efficient.
@@ 4.3 Evaluation Metrics
We evaluate our method on VBench [21], following the protocol in [60, 20]. VBench comprises 16 evaluation dimensions, which are aggregated into three scores: quality, semantic, and total. The quality score is a weighted average of metrics such as subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and dynamic degree, reflecting the technical fidelity of generated videos. The semantic score averages dimensions like object class, multiple objects, human action, color, spatial relationships, and overall consistency, capturing content relevance and semantic coherence. Since we adopt a similar KV cache strategy, our inference runtime matches that of Self Forcing.
@@ 4.4 Main Results
Following prior works [60, 20], we report aggregated scores in Tab. 1, with selected metrics plot in Fig. 2.
@@
. Overall, standard bidirectional teacher models tend to perform better than autoregressive models in many aspects. For instance, the bidirectional model Wan2.1 [54] achieves a total score of 84.26, whereas CausVid [60] attains a lower score of 81.20, likely due to architectural differences and a mismatch between training and inference procedures. Self Forcing [20] attempts to address this issue by incorporating self-rollout during training, which helps reduce the discrepancy and leads to improved results. However, since it is distilled from the teacher using DMD [58], its performance remains largely comparable to that of the bidirectional model with same size.
@@ 4.5 Qualitative Comparison
Here we show the generated videos in Fig. 4 and compare with two other state-of-the-art models including CausVid [60] and Self Forcing [20] and the original Wan2.1 1.3B model. It can be seen that our model is able to generate high quality videos compared to the baseline model. Note that since the texture of our videos are generated with the guidance of an external reward model, the overall style will look different with that of the teacher and baseline models.
Quality of Generated Videos
4.6 Discussion
Our results indicate a useful decoupling between motion learning and appearance refinement in autoregressive video diffusion. The ODE-based initialization provides a compact motion prior, enabling smooth dynamics with only a small set of teacher trajectories, while reward supervision selectively enhances texture without disrupting temporal coherence. This lightweight combination avoids heavy heterogeneous distillation yet yields competitive VBench scores. The findings suggest that strong bidirectional teachers may not be strictly required for high-quality autoregressive generation and that reward-driven refinement offers a promising, data-free alternative for training efficient video generators.
@@ 5.1 Combining Reward with Distillation
Table 2: Performance comparison between our method and one combining both reward and distillation losses. Results are evaluated under similar compute. No further gains observed with more computation.
@@ 6 Conclusion & Future Work
One promising direction for future research is to explore alternative reward models, including both image- and video-based rewards, to further improve the quality of generated videos. Our current approach builds on the findings  [23, 53, 34, 38], which suggest that global structures in images and motions in videos tend to emerge early in the generative process. Directly modeling motion dynamics could potentially further enhance generation quality [48, 4]. In addition, Reward-Instruct [37] shows that high-quality image synthesis can be achieved solely through reward-guided training while bypassing the heavy distillation process. Inspired by this, future work could investigate the possibility of eliminating the initial ODE training phase entirely in favor of a reward-driven learning paradigm for video generation.
@@ 7 Limitations
Although our method demonstrates competitive performances compared to baseline methods, video inherently possesses more dimensions of judgement than images. Our work is purely based on open-sourced reward models, users are recommended to choose the ones that best suit their needs when applying our method to downstream tasks. Our method provides an alternative approach to convert bidirectional teacher models into autoregressive student models. Since standard bidirectional video diffusion models already offer robust guidance, we did not apply our methods to regular models, an avenue we intend to investigate further in future work. Additionally, owing to the limited teacher supervision, we observe some inconsistency in certain videos, which can potentially be addressed with stronger reward models.
Supplementary Material
8 Quality of Generated Videos
@@
9 Full Evaluation Metrics
9.1 More Generated Samples

### v2

@@ 3.3 Baseline Methods
LTX-Video 1.9B $768{\times}512$ 8.98 13.5 80.00 82.30 70.79
Wan2.1 1.3B $832{\times}480$ 0.78 103 84.26 85.30 80.09
@@
NOVA 0.6B $768{\times}480$ 0.88 4.1 80.12 80.39 79.05
Pyramid Flow 2B $640{\times}384$ 6.7 2.5 81.72 84.74 69.62
@@
SkyReels-V2 1.3B $960{\times}540$ 0.49 112 82.67 84.70 74.53
MAGI-1 4.5B $832{\times}480$ 0.19 282 79.18 82.04 67.74
CausVid 1.3B $832{\times}480$ 17.0 0.69 81.20 84.05 69.80
Self Forcing 1.3B $832{\times}480$ 17.0 0.69 84.31 85.07 81.28
@@
3.4 Implementation Details
Similar to Yin et al. (2025); Huang et al. (2025), our model is based on Wan2.1-T2V-1.3B(Wan et al., 2025). The model is first distilled into a 4-step model using DMD. For ODE initialization, we also directly sample 1.4K trajectories from the bidirectional model. Our work uses Self-Rollout during training which is the same as SelfForcing (Huang et al., 2025). For our main method, we only employ loss from the reward models. For our reward plus distillation setting, we train the model using DMD together with reward loss as shown in the previous section. Our method also works for both chunk-wise and frame-wise autoregressive generation. In our implementation, since we only need to sample the ODE trajectory from the teacher model, we don’t need any real training dataset. Thus our overall approach remains data free which is similar to that of Self Forcing (Huang et al., 2025). Similar to previous works (Huang et al., 2025), we enable EMA (Hunter, 1986) during the training process.
Note that our training process eliminates the second distillation stage used in CausVid (Yin et al., 2025) and Self Forcing (Huang et al., 2025). As a result, there is no need to load the teacher and critic models or train the critic model to approximate the generator distribution, making the training process after ODE initialization significantly more lightweight and efficient.
@@ 4.1 Evaluation Metrics
We evaluate our method on VBench (Huang et al., 2024), following the protocol in Yin et al. (2025); Huang et al. (2025). VBench comprises 16 evaluation dimensions, which are aggregated into three scores: quality, semantic, and total. The quality score is a weighted average of metrics such as subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and dynamic degree, reflecting the technical fidelity of generated videos. The semantic score averages dimensions like object class, multiple objects, human action, color, spatial relationships, and overall consistency, capturing content relevance and semantic coherence. Since we adopt a similar KV cache strategy, our inference runtime matches that of Self Forcing. In addition to the aggregated scores, we also report per-dimension results to provide a more fine-grained comparison of visual quality and semantic consistency across different methods.
@@ 4.2 Main Results
Following prior works (Yin et al., 2025; Huang et al., 2025), we report aggregated scores in Tab. 1, with selected metrics plot in Fig. 2.
Overall, standard bidirectional teacher models tend to perform better than autoregressive models in many aspects. For instance, the bidirectional model Wan2.1 (Wan et al., 2025) achieves a total score of 84.26, whereas CausVid (Yin et al., 2025) attains a lower score of 81.20, likely due to architectural differences and a mismatch between training and inference procedures. Self Forcing (Huang et al., 2025) attempts to address this issue by incorporating self-rollout during training, which helps reduce the discrepancy and leads to improved results. However, since it is distilled from the teacher using DMD (Yin et al., 2024a), its performance remains largely comparable to that of the bidirectional model with same size.
@@ 4.3 Qualitative Comparison
Here we show the generated videos in Fig. 3 and compare with two other state-of-the-art models including CausVid (Yin et al., 2025) and Self Forcing (Huang et al., 2025) and the original Wan2.1 1.3B model. It can be seen that our model is able to generate high quality videos compared to the baseline model. Note that since the texture of our videos are generated with the guidance of an external reward model, the overall style will look different with that of the teacher and baseline models.
@@ 5.1 Combining Reward with Distillation
Table 2: Performance comparison between our method and one combining both reward and distillation losses. Results are evaluated under similar compute. No further gains observed with more computation
@@ 6 Conclusion & Future Work
One promising direction for future research is to explore alternative reward models, including both image- and video-based rewards, to further improve the quality of generated videos. Our current approach builds on the findings  (Jiang et al., 2024; Sun et al., 2025; Liu et al., 2025b; Materzyńska et al., 2024), which suggest that global structures in images and motions in videos tend to emerge early in the generative process. Directly modeling motion dynamics could potentially further enhance generation quality (Shi et al., 2024; Chefer et al., 2025). In addition, Reward-Instruct (Luo et al., 2025) shows that high-quality image synthesis can be achieved solely through reward-guided training while bypassing the heavy distillation process. Inspired by this, future work could investigate the possibility of eliminating the initial ODE training phase entirely in favor of a reward-driven learning paradigm for video generation.
@@ 7 Limitations
Although our method demonstrates competitive performance relative to baseline approaches, video generation inherently involves more complex and multi-dimensional judgments than image generation. Our framework relies entirely on open-source reward models, and users may select reward functions that best suit their downstream applications. We propose an alternative pathway for converting bidirectional teacher models into autoregressive student models; however, since bidirectional video diffusion models already provide strong guidance, we do not apply our method to standard non-autoregressive settings, which we leave for future work. Due to limited teacher supervision, we observe occasional inconsistencies in generated videos, which could potentially be mitigated by stronger or more expressive reward models. From a practical standpoint, improving the stability and efficiency of reward-based training for large-scale video models—through better optimization strategies, variance reduction, and scalable reward modeling—remains an important direction for future research.
Appendix A Appendix
A.1 More related works
Motion in Video Diffusion Models
Motion modeling remains a central challenge in video diffusion. Many early video diffusion models, such as Video Diffusion Models (Ho et al., 2022b) primarily focus on spatiotemporal attention mechanisms but treat motion implicitly. More recent efforts like MCDiff (Chen et al., 2023) and VideoJAM (Chefer et al., 2025) attempt to model motion more explicitly, using a flow completion model or joint appearance-motion representations. And Motion-I2V (Shi et al., 2024) further explores motion disentanglement in the image-to-video task, suggesting a broader trend toward treating motion and appearance as separable generative components.
Appendix B Implementation Details
B.1 Computing Infrastructure
B.2 Hyperparameters
Here we list the hyperparameters we used for our experiment. We use a batch size of 8 after training with the ODE trajectory. For the optimizer, we use AdamW with $\beta_{1}=0,\beta_{2}=0.999,\epsilon=1e-8,weight\_decay=0.01$ with a learning rate of 2e62e-6. For EMA, we use a decay of 0.99. When combing distillation loss with image reward loss, we normalize the losses so that they are on the same scale. When comparing with baseline methdos such as DMD or Causvid, we use their author-provided checkpoints.
Appendix C Quality of Generated Videos
@@
Appendix D Generalization to Long Videos
Since our model is directly built upon CausVid (Yin et al., 2025) and Self Forcing (Huang et al., 2025), thus the model can be used to generate long videos as well. E.g. Causvid generates long videos by conditioning on previous generated frames which could be utilized to generate long videos by our method too.
Appendix E Full Evaluation Metrics
@@
E.1 More Generated Samples